1. A political campaign manager wishes to survey a number of voters to estimate



the proportion of those who are in favor of his candidate. If a previous survey



shows that 55% of registered voters plans to vote for his candidate, what is the



minimum sample size required to make his surveys accurate with a 95%



confidence level and a margin of error of 2.5%?

1
Expert's answer
2022-04-27T16:36:16-0400

The formula for error:


"E=z\cdot \sqrt{\cfrac{\hat{p}\cdot (1-\hat{p})}{n}}."


Here

"E\ -" the error, "E=0.025" ;

"z\ -" z-score, for 95% confidence level "z=1.96" ;

"\hat{p}\ -" the sample proportion, "\hat{p}=0.55;"

"n\ -" the sought sample size.


So,


"n=\cfrac{z^2\cdot\hat{p}\cdot (1-\hat{p})}{E^2}"

"=\cfrac{1.96^2\cdot0.55\cdot (1-0.55)}{0.025^2}=1521.3."

The minimum sample size is 1522.